OpenStudy (crashonce):

2 Litres of a liquid containing 60% of alcohol. How much alchohol would be added to obtain one containing 75% alcohol

OpenStudy (crashonce):

@ganeshie8 asking again to try to get easier understanding

OpenStudy (kainui):

Is this 75% by total content or by volume?

OpenStudy (kainui):

(is this a math or a chemistry problem?)

OpenStudy (mathmath333):

two litres contain \(60\%\) alcohol that means \(1.2\) litres is alcohol and \(0.8\) litres is other liquid let the \(x\) litre alcohol should be added to make it \(75\%\) of the liquid so \(\large\tt \begin{align} \color{black}{(1.2+x)=(75\%)(2)\\~\\ (1.2+x)=(\dfrac{75}{100})(2)\\~\\ }\end{align}\)

OpenStudy (mathmath333):

this is an easy problem of alligation and mixture

OpenStudy (crashonce):

what is ur answer

OpenStudy (mathmath333):

wait sry \(\large\tt \begin{align} \color{black}{(1.2+x)=(75\%)(2+x)\\~\\ (1.2+x)=(\dfrac{75}{100})(2+x)\\~\\ }\end{align}\)

OpenStudy (mathmath333):

this should be correct

OpenStudy (crashonce):

wait so what is ur answer

OpenStudy (mathmath333):

\(2+1.2=3.2\\ 3.2\times 0.75=2.4=1.2+1.2\)

OpenStudy (mathmath333):

when i started studying these types of problem , they were so confusing